An investor is forming a portfolio by holding $8,000 in stock A, which has a beta of 1.50, and $5,000 in stock B, which has a beta of 0.92. Find portfolio beta.
Answer #1

Total investment=(8000+5000)=$13000

Portfolio beta=Respective betas*Respective investment weights

=(8000/13000*1.5)+(5000/13000*0.92)

which is equal to

=1.28(Approx).